OpenStudy (anonymous):

what is the nth term of sequence 4,9,16,25,.......

OpenStudy (raden):

hint : 4 = (1+1)^2 9 = (2+1)^2 16 = (3+1)^2 25 = (4+1)^2 36 = (5+1)^2 so on ....